OpenStudy (anonymous):

In 1986 there were 1657 Bus drivers.In 2007 the numbers had dropped to 1048.Assume that the decline in the number of bus drivers generates an arithmetic sequence. Determine the decline of bus driver each year. I was just introduced to the formula. tn=t1+(n-1)d. I would like to solve problem using that formula if possible. Thanks

OpenStudy (kropot72):

The number of years between 1986 and 2007 = 2007 - 1986 = 21 years therefore the number of terms in the arithmetic progression = 21 = n The decline of bus drivers each year is the common difference = d \[n ^{th}\ term=1^{st}\ term+(n-1)d\] \[1048=1657+(21-1)d\] \[d=\frac{1048-1657}{21-1}=you\ can\ calculate\]

OpenStudy (anonymous):

only think I dont get is why 1048 is were it is

OpenStudy (anonymous):

I understand everything else but Im not following that

OpenStudy (kropot72):

\[1048=n ^{th}\ term=last\ term=total\ drivers\ 2007\]

OpenStudy (anonymous):

ok thanks

OpenStudy (kropot72):

You're welcome :)